OpenStudy (anonymous):

A patient is to take 4 1/3 tablespoons of medicine per day in 4 equally doses. How much medcine is to be taken in each dose? Any help?

jimthompson5910 (jim_thompson5910):

To find the answer, divide 4 1/3 by 4 to get \[\large \frac{4\frac{1}{3}}{4}\] \[\large \frac{4+\frac{1}{3}}{4}\] \[\large \frac{\frac{12}{3}+\frac{1}{3}}{4}\] \[\large \frac{\frac{12+1}{3}}{4}\] \[\large \frac{\frac{13}{3}}{4}\] \[\large \frac{\frac{13}{3}}{\frac{4}{1}}\] \[\large \frac{13}{3}\frac{1}{4}\] \[\large \frac{13*1}{3*4}\] \[\large \frac{13}{12}\] \[\large \frac{12+1}{12}\] \[\large \frac{12}{12}+\frac{1}{12}\] \[\large 1+\frac{1}{12}\] \[\large 1\frac{1}{12}\] So \[\large \frac{4\frac{1}{3}}{4}=1\frac{1}{12}\] So each dose contains \[1\frac{1}{12}\] tablespoons of medicine.

OpenStudy (anonymous):

much apperciated